. Given the total utility function (T u) = 150x+40x2 = x 3 then Derive marginal utility function and find the value of x at which total utility is maximum 


1
Expert's answer
2021-04-13T07:16:33-0400

"MU = TU'(x) = 150 + 80x - 3x^2."

Total utility is maximized, if MU = 0, so:

"150 + 80x - 3x^2 = 0,"

D = 8,200,

"x1 = \frac{-80 + 8,200^{0.5}} {2×3} = 1.76."

x2 < 0, so is not suitable for our case.